OpenStudy (anonymous):

Find the slope of the line through two points:(2,-2)(5,7)

OpenStudy (anonymous):

3

OpenStudy (anonymous):

7+2/5-2 = 3

OpenStudy (anonymous):

y2-y1/(x2-x1)= 7-(-2)/(5-2)=9/3=3

OpenStudy (anonymous):

So if then it wants in slope intercept form? y=3x + 7?